OpenStudy (anonymous):

sqrt{5a}+5sqrt{45a^3} sqrt{5a}+5sqrt{45a^3} @Mathematics

OpenStudy (anonymous):

\[\sqrt{5a}+5\sqrt{45a ^{3}}\]

jhonyy9 (jhonyy9):

sqrt5a +5sqrt(9*5*a2*a)=sqrt5a+5*3*asqrt(5a)=sqrt5a+15asqrt5a= =sqrt5a(1+15a)